32.0 g of copper pellets are removed from a 300° C oven and immediately dropped into 110 mL of water at 22.0 ° C in an insulated cup What will the new water temperature be? Express your answer with the appropriate units You may want to review (Pages 530-531) Value Units
Answer #1

here,

mass of copper , m1 = 32 g = 0.032 kg

mass of water , m2 = 110 mL = 110 g = 0.11 kg

let the new temperature be Tf

using conservation of heat energy

m1 * Cc * ( 300 - Tf) = m2 * Cw * ( Tf - 22)

0.032 * 385 * ( 300 - Tf ) = 0.11 * 4186 * ( Tf - 22)

solving for Tf

Tf = 29.2 degree

the final temperature is 29.2 degree